OpenStudy (anonymous):

Which of the following pairs of equations represent parallel lines? A. y = 2x + 4 and y = -1/2x + 4 B. y = 3x - 14 and y = 3x + 2 C. y = 7x - 5 and y = -7x + 5 D. y = 4x + 2 and y = 1/4x - 2

OpenStudy (anonymous):

@jojo4eva @Preetha

OpenStudy (anonymous):

Parallel lines have the same slope..so which ones have the same slope?

OpenStudy (anonymous):

All of them are in the form of: y = mx + b Where M is slope.

OpenStudy (anonymous):

So which one's have the same 'm'?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

y = 2x + 4 and y = -1/2x + 4 ^ Slope ^ Slope The slope of the first one is 2, and the slope of the 2nd one is -1/2. 2 and -1/2 are NOT the same, so they are not parallel lines.

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

No..that's close but notice that the first one has 7, and the other one has -7.

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yep, B is correct, both of them have a slope of 3.

OpenStudy (jojo4eva):

the answer is B

OpenStudy (anonymous):

THANK YOU @iGreen. and you are late @jojo4eva

OpenStudy (jojo4eva):

i kno my internet was acting up but sorry im good now

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Okay so could you help me with a few more ? @jojo4eva

OpenStudy (jojo4eva):

ya can you givr me a moment tho?? ill be right back

OpenStudy (anonymous):

ok

OpenStudy (jojo4eva):

ok @tiffany__ im ready when you are